# Rank-DistiLLM: Closing the Effectiveness Gap Between Cross-Encoders and LLMs for Passage Re-Ranking

### v3

@@ 4.1 In-Domain Effectiveness
Table 2: Effectiveness in nDCG@10 micro-averaged across all queries from a collection from the TIREx framework [32]. The macro-averaged geometric mean is computed across all corpora. The highest and second-highest scores per corpus are bold and underlined, respectively. $\dagger$ denotes a statistically significant difference from the underlined monoELECTRA model ($p<0.05$, Holm-Bonferroni-corrected). The dataset used to fine-tune each model is provided for context (RDL: our Rank-DistiLLM dataset).
@@
First Stage .510$\dagger$ .405 .173 .364$\dagger$ .586$\dagger$ .012 .424$\dagger$ .332$\dagger$ .467$\dagger$ .374 .268$\dagger$ .447$\dagger$ .364$\dagger$ .290
RankZephyr RGPT+ .528$\dagger$ .364 .203 .303 .767$\dagger$ .009 .542$\dagger$ .423$\dagger$ .560 .453$\dagger$ .299 .512 .508$\dagger$ .321
monoT53B MSM .537$\dagger$ .392 .181 .279 .603$\dagger$ .011 .545$\dagger$ .407$\dagger$ .514 .375 .308$\dagger$ .458$\dagger$ .476 .305
RankT53B MSM .592 .421$\dagger$ .215 .336$\dagger$ .713 .010 .538$\dagger$ .415$\dagger$ .528 .406 .307$\dagger$ .459$\dagger$ .468 .323
mELECBase MSM .512$\dagger$ .326$\dagger$ .155$\dagger$ .252$\dagger$ .667 .008 .436$\dagger$ .337$\dagger$ .491$\dagger$ .364$\dagger$ .255$\dagger$ .456$\dagger$ .406 .271
mELECBase TWOLAR .570$\dagger$ .305$\dagger$ .180 .292 .653$\dagger$ .009 .486$\dagger$ .354$\dagger$ .523 .404 .267$\dagger$ .519 .434 .292
@@
.570$\dagger$ .369 .210 .313$\dagger$ .716 .008 .546$\dagger$ .424$\dagger$ .572$\dagger$ .417 .301$\dagger$ .526 .504$\dagger$ .320
@@ 4.4 Listwise Fine-Tuning
Our newly proposed listwise ADR-MSE loss function produces a marginally less effective model at 0.002 (single-stage) and 0.001 (two-stage) lower nDCG@10 compared to using RankNet averaged across TREC DL 2019 / 2020 and BM25 / ColBERTv2 for initial retrieval. Since the effectiveness is practically equal and monoELECTRALarge fine-tuned using RankNet already reaches the effectiveness of RankZephyr, we conclude that listwise loss functions are (currently) unnecessary for distillation from LLMs. More complex ranking tasks may benefit from listwise losses, warranting further research for confirmation.

### v4

@@ 4.1 In-Domain Effectiveness
Table 2: Effectiveness in nDCG@10 micro-averaged across all queries from a collection from the TIREx framework [32]. The macro-averaged geometric mean is computed across all corpora. The highest and second-highest scores per corpus are bold and underlined, respectively. $\dagger$ denotes a statistically significant difference from the underlined monoELECTRA model ($p<0.05$, Holm-Bonferroni-corrected). The dataset used to fine-tune each model is provided for context (RDL: our Rank-DistiLLM dataset). [^5] This table very slightly differs from the table in the official proceedings: it correctly reports micro-averaged results and the geometric mean.
@@
First Stage .516$\dagger$ .405 .177 .364$\dagger$ .586$\dagger$ .012 .424$\dagger$ .259$\dagger$ .467$\dagger$ .385 .281$\dagger$ .447$\dagger$ .364$\dagger$ .286
RankZephyr RGPT+ .534$\dagger$ .364 .213 .303 .767$\dagger$ .009 .542$\dagger$ .349$\dagger$ .560 .460$\dagger$ .314 .512 .508$\dagger$ .320
monoT53B MSM .590 .415$\dagger$ .188 .323$\dagger$ .649 .011 .526$\dagger$ .345$\dagger$ .529 .395 .319$\dagger$ .474$\dagger$ .469 .313
RankT53B MSM .598 .421$\dagger$ .227 .336$\dagger$ .713 .010 .538$\dagger$ .353$\dagger$ .528 .406 .323$\dagger$ .459$\dagger$ .468 .322
mELECBase MSM .517$\dagger$ .326$\dagger$ .164$\dagger$ .252$\dagger$ .667 .008 .436$\dagger$ .254$\dagger$ .491$\dagger$ .366$\dagger$ .268$\dagger$ .456$\dagger$ .406 .267
mELECBase TWOLAR .576$\dagger$ .305$\dagger$ .186 .292 .653$\dagger$ .009 .486$\dagger$ .263$\dagger$ .523 .406 .281$\dagger$ .519 .434 .288
@@
.575$\dagger$ .369 .221 .313$\dagger$ .716 .008 .546$\dagger$ .344$\dagger$ .572$\dagger$ .419 .316$\dagger$ .526 .504$\dagger$ .318
@@ 4.4 Listwise Fine-Tuning
Our newly proposed listwise ADR-MSE loss function produces a marginally less effective model at 0.002 (single-stage) and 0.001 (two-stage) lower nDCG@10 compared to using RankNet averaged across TREC DL 2019 / 2020 and BM25 / ColBERTv2 for initial retrieval. Since the effectiveness is practically equal, we conclude that listwise loss functions are (currently) unnecessary for distillation from LLMs. Future more complex tasks may benefit from listwise distillation.